# NMPCB: A Lightweight and Safety-Critical Motion Control Framework

### v1

@@ I Introduction
In the design of robot motion planners and controllers, safety has always been the paramount requirement, considering the rapid development of robot systems[1]. The control method based on CBF has garnered significant attention for offering a simple yet safety-guaranteed approach to robot motion control[2]. In particular, the control method that integrates CBF and MPC have demonstrated superior performance in safety control by predicting future states [3]. However, this method often fails to obtain an optimal solution within a short time, especially for nonlinear kinematic systems where solution failure frequently occurs. Consequently, it also demands a high level of precision from the reference trajectory.
This letter presents a novel framework that integrates a neural network-based planner with an improved CBF-based MPC controller to ensure safety and real-time performance in robot motion control. The proposed framework serves both as a real-time local planner and as a controller.
@@ I-A2 Optimization-Based Controller
The CBF serves as a safety boundary function for the control process, providing assurance of the robot’s safety[2][14]. CBF is particularly well-suited for collision avoidance constraints and have numerous applications in the fields of robotic manipulators [15] and vehicular systems [16]. In discrete-time systems, discrete-time control barrier function (DCBF) shows good performance in terms of safety constraints [17]. Research has been conducted on DCBF in the control and planning problems of multi-step optimization [18].
@@ I-B Contributions
  • We propose an encoder-decoder path planning architecture which can determine the subsequent target point using historical path information and employs the Dubins curve to the target point as the reference path input to the control module.
  • Based on our planning architecture, we have improved the MPC-DCBF formulation to better accommodate nonlinear kinematic constraints and ensure its real-time performance.
  • We conducted numerical simulations and real-world experiments for the proposed framework in this letter, and performed comparative experiments with several baseline schemes.
@@ II-A Model Predictive Control
If the distance to obstacles is employed as a collision avoidance constraint, then the MPC control formulation (MPC-DC) at time $t$ is given as follows:
@@
$\quad g(\mathbf{x}_{t+k|t})\geq 0,\,k=0,\ldots,N-1.$ (2e)
Here, The functions $p(\cdot)$ and $q(\cdot)$ in (2a) represent the terminal cost and the cost associated with tracking the reference trajectory, respectively. Constraint (2d) indicates that the initial state of the robot is $\mathbf{x}_{t}$ and the subsequent states are predicted via the system dynamics given in (2b), and constraint (2d) also specifies the terminal constraint. Both state and input constraints are provided by (2c). The collision avoidance constraint is given by (2e), which can be defined under various Euclidean norms.
We can also regard the MPC-DC as a path planning module. The path planning problem can be formulated as planning a collision-free trajectory for the robot from the initial state $\mathbf{x}_{0}$ to the target state $\mathbf{x}_{goal}$. If we set the value of $k$ in the optimization model (2) to 0 and solve for the optimal state $\mathbf{x}$, this becomes a standard path planning problem. We will employ an encoder-decoder model to address the path planning problem, the specifics of which will be discussed in Section III-A.
@@ II-B Control Barrier Function
$0\leq\gamma(\mathbf{x})<1,\forall\mathbf{x}\in\mathcal{S},\exists\mathbf{u}\in\mathcal{U},$
Let $\gamma_{k}:=\gamma(\mathbf{x}_{k})$. Satisfying (3) implies $h(f(\mathbf{x},\mathbf{u}))\geq\gamma_{K}h(\mathbf{x})$, that is, the lower bound of the DCBF decreases exponentially with the decay rate $\gamma_{k}$ [18].
@@
The formulation proposed later in [nmpc dcbf 33] introduces a slack variable $\omega$ to balance feasibility and safety as follows:
@@
The MPC control formulation based on DCBF constraints (MPC-DCBF) at time $t$ is presented as follows:
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1,$ (7d)
@@
The optimization formulation (7) is fundamentally similar to the optimization formulation (2). In formulation (7), $U=[\mathbf{u}^{T}_{t|t},\ldots,\mathbf{u}^{T}_{t+N-1|t}]$ and $\Omega=[\omega_{0},\ldots,\omega_{N_{CBF-1}}]$ are the joint input and relaxation variables, respectively. $\psi$ is the penalty function for the relaxation variable.
Figure 2: Figure illustrating the Neural Dubins Model.
III NMPCB Motion Control Framework
In this section, we introduce the NMPCB framework, which consists mainly of two components: neural network-based path planner and the DCBF-based MPC controller.
III-A Neural Dubins Model
In this section, we describe our model formulation. As shown in Fig. 2 and Fig. 3, our model mainly consists of three components: the reference path encoder, the obstacle encoder, and the decoder.
The reference path encoder receives information from the start point $\mathbf{x}_{0}=[x_{0},y_{0},\theta_{0}]$ and the goal point $\mathbf{x}_{goal}=[x_{goal},y_{goal},\theta_{goal}]$ as input, where $x$ and $y$ represent the Cartesian coordinates, and $\theta$ denotes the heading angle. Within the model, a Dubins curve[23] is first generated from the $\mathbf{x}_{0}$ to $\mathbf{x}_{goal}$, and the coordinates of each point $P_{k}=[x_{k},y_{k},\theta_{k}]$ on the curve along are sequentially input into a single-layer LSTM module to obtain the hidden state $H$. LSTM networks exhibit superior adaptability to motion continuity and enhanced robustness to noisy or irregular path data when processing path sequences.
The obstacle encoder receives obstacle information as input. Each obstacle can be represented as $O_{i}=[x_{i},y_{i},r_{i}]$, where $x$ and $y$ represent the center, and $r$ denotes the obstacle radius. The model maps the obstacle information onto a $50\times 50$ map matrix, where matrix values are set to $1$ at grid points with obstacles and 00 at obstacle-free points, thus, the dimension of each map matrix is $(50,50,1)$. This map matrix is then fed into a hierarchical feature extraction architecture comprising two convolutional downsampling modules (CDM) and a fully connected compression module, which ultimately outputs a low-dimensional feature vector $Z$. Each convolutional downsampling module comprises individual two-dimensional convolutional layers with a kernel size of $3\times 3$. Each convolutional layer is connected with batch normalization (BN), an activation function, the rectified linear unit (ReLU), and a downsampling layer, max pooling, with filters of size $2\times 2$.
Figure 3: Figure illustrating the CDM.
Subsequently, the hidden state $H$ of the reference path encoder, the feature vector $Z$ of the obstacle encoder, the endpoint $\mathbf{x}_{goal}$, and all path points from the start point to the current point $\mathbf{x}_{0:t}$ are input into the decoder. It then outputs the future states $\hat{\mathbf{{x}}}_{t+N_{pred}}$ within a specified time horizon $N_{pred}$,
$\mathcal{D}(H,Z,\mathbf{x}_{goal},\mathbf{x}_{0:t})\rightarrow\hat{\mathbf{{x}}}_{t+N_{pred}}.$ (8)
The decoder network $\mathcal{D}$ is composed of a single-layer LSTM network followed by a fully connected layer. We use a mean-squared-error (MSE) loss between the predicted state $\hat{\mathbf{{x}}}_{t+N_{pred}}$ and the label next state $\mathbf{{x}}_{t+N_{pred}}$ during training, as follows:
$loss=\frac{1}{N}\sum_{k=0}^{N}\left\|\hat{\mathbf{x}}_{k+N_{pred}}-\mathbf{x}_{k+N_{pred}}\right\|^{2}.$ (9)
Here, $N$ denotes the total number of planning steps.
After deriving the predicted state $\hat{\mathbf{{x}}}_{t+N_{pred}}$, a Dubins curve is constructed from the current point $\mathbf{x}_{t}$ to $\hat{\mathbf{{x}}}_{t+N_{pred}}$ to serve as a reference trajectory that is transmitted to the control module.
III-B Dual DCBF Constraint
@@
$\mathcal{O}_{i}:=\{y\in\mathbb{R}^{l}:A^{\mathcal{O}_{i}}y\leq b^{\mathcal{O}_{i}}\}$ (10)
where $b^{\mathcal{O}_{i}}\in R^{s^{\mathcal{O}_{i}}}$. $s^{\mathcal{O}_{i}}$ represent the number of facets of polytopic sets for the $i$-th obstacle.
Let the state of the robot be $\mathbf{x}\in\mathbb{R}^{n}$ with its discrete-time dynamics as defined in (1) , $\mathcal{R}(\mathbf{x})$ represent the center of the robot. For $\mathbf{x}\in\mathcal{X}$, the computation of the squared minimum distance between the $i$-th obstacle $\mathcal{O}_{i}$ and the center $\mathcal{R}(\mathbf{x})$ can be be denoted by the function $h_{i}(\mathbf{x})$ formula:
@@
[…]
It can be noted that Equation (11) formulates a quadratic programming (QP) problem, thus representing a convex optimization problem. To ensure safety throughout the motion process, we apply the same DCBF constraint to each obstacle in relation to the robot. Thus, the safety set is defined as:
@@
We note that in the DCBF constraint (4), the computation of $h(f(\mathbf{x},\mathbf{u}))$ is required, which leads to the presence of non-differentiable implicit constraints in the optimization formulation and can consume significant computational time. Next, we introduce the dual DCBF constraints to address this issue.
@@
s.t. $\quad\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1$ (13b)
$\quad\lambda^{\mathcal{O}_{i}}\geq 0$ (13c)
Here $\lambda A^{\mathcal{O}_{i}}$ is the normal vector of the separating hyperplane.
@@
Then, we can transform the DCBF constraints into a more stronger form:
@@
This represents a stronger DCBF constraint that satisfies the (4). The substitution of the DCBF constraint with (16) requires satisfying (13b) and (13c), as follows:
@@
According to the Strong Duality Theorem (14), $\lambda^{\mathcal{O}_{i}*}$ satisfying (10b)-(10c) such that for all $x\in\mathcal{X}$
$\bar{g}_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}*})=(A^{\mathcal{O}_{i}*}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}*})^{T}\lambda^{\mathcal{O}_{i}*}=h_{i}(\mathbf{x}).$ (18)
@@
Similarly, we can apply the aforementioned method to optimization problem (11) to establish stronger DCBF constraints. Let $\mathbf{y}^{\mathcal{O}_{i}}$ be any feasible solution to (11). Since (11) is a minimization problem, we can derive the following inequality relationship:
$\bar{h}_{i}(\mathbf{x},y^{\mathcal{O}_{i}},y^{\mathcal{R}}):=\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}\geq h_{i}(\mathbf{x}).$ (19)
@@
Comparing with equation (4), we can observe the transformation of the DCBF constraints:
@@
Integrating the analysis presented above, we define the dual DCBF constraints as follows:
@@ III-C Optimization Formulation
We apply the dual DCBF constraints to the MPC algorithm as collision avoidance constraints to construct a multi-step optimization model, thereby providing enhanced assurance of safety in motion control. The optimization formulation for MPC-DUAL-DCBF (MDD) at time $t$ is presented as follows:
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23e)
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23f)
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{CBF}-1$ (23g)
In the formulation, $y_{t}^{\mathcal{O}_{i}*}$ represents the optimal solution obtained by precomputing the minimum distance at time $t$ through (11). Constraint (23e) represents the dual DCBF constraint, while (23f)-(23g) denote the feasibility conditions of the dual DCBF constraint. The optimization formulation (23) illustrates only the dual DCBF constraint between obstacle $O_{i}$ and the region $R$, however, during the motion control process, the corresponding dual DCBF constraints are applied to each pair of obstacles and the robot.
If the same dual DCBF constraints are imposed at each time step within a multi-step optimization model, it would result in a substantial computational burden. To reduce this complexity, we roll out the time $k$ and replace the RHS of each DCBF constraint with $h_{i}(\mathbf{x}_{t})$, as follows:
@@
Replacing (23e) with Equation (24) can accelerate the computation. This modification affects neither the system feasibility nor its safety.

### v2

@@ Abstract
Index Terms:
Integrated planning and control, integrated planning and learning, collision avoidance, optimization and optimal control.
@@ I Introduction
In the design of robot motion planners and controllers, safety has always been the paramount requirement, considering the rapid development of robot systems[1]. The control method based on CBF has garnered significant attention for offering a simple yet safety-guaranteeing approach to robot motion control[2]. In particular, the control method that integrates CBF and MPC has demonstrated superior performance in safety control by predicting future states [3]. However, this method often fails to obtain an optimal solution within a short computation time, especially for nonlinear kinematic systems where solution failure frequently occurs. Consequently, it also demands a high level of precision from the reference trajectory.
This paper presents a novel framework that integrates a neural network-based planner with an improved CBF-based MPC controller to ensure safety and real-time performance in robot motion control. The proposed framework serves both as a real-time local planner and as a controller.
@@ I-A2 Optimization-Based Controller
The CBF serves as a safety boundary function for the control process, providing assurance of the robot’s safety[2][14]. CBF is particularly well-suited for collision avoidance constraints and has numerous applications in the fields of robotic manipulators [15] and vehicular systems [16]. In discrete-time systems, discrete-time control barrier function (DCBF) shows good performance in terms of safety constraints [17]. Research has been conducted on DCBF in the control and planning problems of multi-step optimization [18].
@@ I-B Contributions
  • We propose an encoder-decoder path planner that can predict a collision-free next target point using only historical trajectories and obstacle information, while significantly reducing the time overhead of the planner.
  • We improve the MPC-DCBF formulation by constructing the dual form of obstacle avoidance constraints, which converts the implicit CBF constraints into explicit ones, thereby allowing it to better adapt to obstacle avoidance tasks and ensuring real-time performance.
  • We conduct numerical simulations and real-world experiments for the proposed framework in this paper, and perform comparative experiments with several baseline schemes.
@@ II-A Model Predictive Control
If the distance to obstacles is employed as a collision avoidance constraint, then the MPC control formulation at time $t$ is given as follows:
@@
$\quad g(\mathbf{x}_{t+k+1|t})\geq 0,\,k=0,\ldots,N-1.$ (2e)
Here, the functions $p(\cdot)$ and $q(\cdot)$ in (2a) represent the terminal cost and the cost associated with tracking the reference trajectory, respectively. Constraint (2d) indicates that the initial state of the robot is $\mathbf{x}_{t}$ and the subsequent states are predicted via the system dynamics given in (2b), and constraint (2d) also specifies the terminal constraint. Both state and control input constraints are provided by (2c). The collision avoidance constraint is given by (2e), which can be defined using various Euclidean norms. $N$ denotes the prediction horizons.
We can also regard the MPC as a path planning module. The path planning problem can be formulated as planning a collision-free trajectory for the robot from the initial state $\mathbf{x}_{0}$ to the target state $\mathbf{x}_{\text{goal}}$. If we set the length of the prediction horizons $N$ in the optimization model (2) to $1$ and solve for the optimal state $\mathbf{x}$, this becomes a standard path planning problem. We will employ an encoder-decoder model to address the path planning problem, the specifics of which will be discussed in Section III-A.
@@ II-B Control Barrier Function
$0\leq\gamma(\mathbf{x})<1,\forall\mathbf{x}\in\mathcal{S},\ \exists\mathbf{u}\in\mathcal{U}.$
Let $\gamma_{k}:=\gamma(\mathbf{x}_{k})$. Satisfying (4) implies $h(f(\mathbf{x},\mathbf{u}))\geq\gamma_{k}h(\mathbf{x})$, that is, the lower bound of the DCBF decreases exponentially with the decay rate $\gamma_{k}$ [18]. We denote $\mathcal{K}(\mathbf{x})$ as
@@
The formulation proposed later in [18] introduces a slack variable $\omega$ to balance feasibility and safety as follows:
@@
The MPC control formulation based on DCBF constraint (MPC-DCBF) at time $t$ is presented as follows:
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1,$ (7d)
@@
Figure 2: Illustration of the Neural Dubins Model.
The optimization formulation (7) is fundamentally similar to the optimization formulation (2). In formulation (7), $U=[\mathbf{u}^{T}_{t|t},\ldots,\mathbf{u}^{T}_{t+N-1|t}]$ and $\Omega=[\omega_{0},\ldots,\omega_{N_{\text{CBF}-1}}]$ are the control input and relaxation variables, respectively. $\psi$ is the penalty function for the relaxation variable. $N_{\text{CBF}}$ denotes the safety horizon, which must satisfy $N_{\text{CBF}}\leq N$.
II-C Obstacle Avoidance Problem
@@
$\mathcal{O}_{i}:=\{y\in\mathbb{R}^{l}:A^{\mathcal{O}_{i}}y\leq b^{\mathcal{O}_{i}}\},$ (8)
where $b^{\mathcal{O}_{i}}\in R^{s^{\mathcal{O}_{i}}}$. $s^{\mathcal{O}_{i}}$ represents the number of facets of polytopic sets for the $i$-th obstacle.
Let the state of the robot be $\mathbf{x}\in\mathbb{R}^{n}$ with its discrete-time dynamics as defined in (1) , and let $\mathcal{R}(\mathbf{x})$ represent the center of the robot. For $\mathbf{x}\in\mathcal{X}$, the computation of the squared minimum distance between the $i$-th obstacle $\mathcal{O}_{i}$ and the center $\mathcal{R}(\mathbf{x})$ can be denoted by the function $h_{i}(\mathbf{x})$ formula:
@@
$\quad\quad\text{s.t.}\quad\quad A^{\mathcal{O}_{i}}y^{\mathcal{O}_{i}}\leq b^{\mathcal{O}_{i}}.$ (9b)
It can be noted that (9) formulates a quadratic programming (QP) problem, thus representing a convex optimization problem. To ensure safety throughout the motion process, we apply the same DCBF constraint to each obstacle in relation to the robot. Thus, the safety set is defined as:
@@
III NMPCB Motion Control Framework
In this section, we introduce the NMPCB framework, which consists mainly of two components: the neural network-based path planner and the DCBF-based MPC controller.
III-A Neural Dubins Model
In this section, we describe our encoder-decoder model formulation. As shown in Fig. 2 and Fig. 3, our model mainly consists of three components: the reference path encoder, the obstacle encoder, and the decoder.
The reference path encoder receives information from the start point $\mathbf{x}_{0}=[x_{0},y_{0},\theta_{0}]$ and the goal point $\mathbf{x}_{\text{goal}}=[x_{\text{goal}},y_{\text{goal}},\theta_{\text{goal}}]$ as input, where $x$ and $y$ represent the position of the robot, and $\theta$ denotes the heading angle of the robot. Within the model, a Dubins curve[23] is first generated from $\mathbf{x}_{0}$ to $\mathbf{x}_{\text{goal}}$, and the coordinates of each point $P_{k}=[x_{k},y_{k},\theta_{k}]$ on the curve are sequentially input into a single-layer LSTM [24] module to obtain the hidden state $H$. LSTM networks exhibit superior adaptability to motion continuity and enhanced robustness to noisy or irregular path data when processing path sequences.
The obstacle encoder receives obstacle information as input. Each obstacle can be represented as $O_{i}=[x_{i},y_{i},r_{i}]$, where $x$ and $y$ represent the center, and $r$ denotes the obstacle radius. The model maps the obstacle information onto a $50\times 50$ map matrix, where matrix values are set to $1$ at grid points with obstacles and 0 at obstacle-free points, thus, the dimension of each map matrix is $(50,50,1)$. This map matrix is then fed into a hierarchical feature extraction architecture comprising two convolutional downsampling modules (CDM) and a fully connected compression module, which ultimately outputs a low-dimensional feature vector $Z$. Each CDM comprises individual two-dimensional convolutional layers with a kernel size of $3\times 3$. Each convolutional layer is followed by batch normalization (BN), ReLU activation, and a $2\times 2$ max pooling layer for downsampling.
Figure 3: Illustration of the CDM.
Subsequently, the hidden state $H$ of the reference path encoder, the feature vector $Z$ of the obstacle encoder, the endpoint $\mathbf{x}_{\text{goal}}$, and all path points from the start point to the current point $\mathbf{x}_{0:t}$ are input into the decoder. It then outputs the future states $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$ within a specified time horizon $N_{\text{pred}}$,
$\mathcal{D}(H,Z,\mathbf{x}_{\text{goal}},\mathbf{x}_{0:t})\rightarrow\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}.$ (11)
The decoder network $\mathcal{D}$ is composed of a single-layer LSTM network followed by a fully connected layer. We use a mean squared error (MSE) loss between the predicted state $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$ and the label next state $\mathbf{{x}}_{t+N_{\text{pred}}}$ during training. The loss function $\mathcal{L}$ is defined as follows:
$\mathcal{L}(\hat{\mathbf{x}},\mathbf{x})=\frac{1}{N}\sum_{k=0}^{N}\left\|\hat{\mathbf{x}}_{k+N_{\text{pred}}}-\mathbf{x}_{k+N_{\text{pred}}}\right\|^{2}_{2}.$ (12)
Here, $N$ denotes the total number of planning steps.
After deriving the predicted state $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$, a Dubins curve is constructed from the current point $\mathbf{x}_{t}$ to $\hat{\mathbf{{x}}}_{t+N_{\text{pred}}}$ to serve as a reference trajectory that is transmitted to the control module. We refer to the above-mentioned neural network-based path planner as the Neural Dubins Model.
III-B Dual DCBF Constraint
We note that in the DCBF constraint (4), the computation of $h(f(\mathbf{x},\mathbf{u}))$ is required, which leads to the presence of non-differentiable implicit constraints in the optimization formulation and can consume significant computational time. Next, we introduce the dual DCBF constraint to address this issue.
@@ III-B Dual DCBF Constraint
s.t. $\quad\|\lambda^{\mathcal{O}_{i}}A^{\mathcal{O}_{i}}\|_{2}\leq 1,$ (13b)
$\quad\lambda^{\mathcal{O}_{i}}\geq 0.$ (13c)
Here $(A^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}}$ is the normal vector of the separating hyperplane.
@@
Then, we can transform the DCBF constraint into a stronger form:
@@
This represents a stronger DCBF constraint that satisfies the DCBF constraint (4). The substitution of the DCBF constraint with (16) requires satisfying (13b) and (13c), as follows:
@@
According to the Strong Duality Theorem (14), $\exists\ \lambda^{\mathcal{O}_{i}*}$ satisfying (13b) and (13c) such that for all $x\in\mathcal{X}$,
$\bar{g}_{i}(\mathbf{x},\lambda^{\mathcal{O}_{i}*})=(A^{\mathcal{O}_{i}}\mathcal{R}(\mathbf{x})-b^{\mathcal{O}_{i}})^{T}\lambda^{\mathcal{O}_{i}*}=h_{i}(\mathbf{x}).$ (18)
@@
Similarly, we can apply the aforementioned method to optimization problem (9) to establish stronger DCBF constraint. Let $\mathbf{y}^{\mathcal{O}_{i}}$ be any feasible solution to (9). Since (9) is a minimization problem, we can derive the following inequality relationship:
$\bar{h}_{i}(\mathbf{x},y^{\mathcal{O}_{i}}):=\|y^{\mathcal{O}_{i}}-\mathcal{R}(\mathbf{x})\|_{2}^{2}\geq h_{i}(\mathbf{x}).$ (19)
@@
Comparing with (4), we can observe the transformation of the DCBF constraint:
@@
Integrating the analysis presented above, we define the dual DCBF constraint as follows:
@@ III-C Optimization Formulation
We apply the dual DCBF constraint to the MPC algorithm as collision avoidance constraints to construct a multi-step optimization model, thereby providing enhanced assurance of safety in motion control. The optimization formulation for MPC-DUAL-DCBF (MDD) at time $t$ is presented as follows:
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1$ (23e)
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1$ (23f)
@@
$\quad\quad\text{for}\ k=0,\ldots,N_{\text{CBF}}-1$ (23g)
In the formulation, $y_{t}^{\mathcal{O}_{i}*}$ represents the optimal solution obtained by precomputing the minimum distance at time $t$ through (9). Constraint (23e) represents the dual DCBF constraint, while (23f)-(23g) denote the feasibility conditions of the dual DCBF constraint. The optimization formulation (23) illustrates only the dual DCBF constraint between obstacle $\mathcal{O}_{i}$ and the robot $\mathcal{R}$, however, during the motion control process, the corresponding dual DCBF constraint are applied to each pair of obstacles and the robot.
If the same dual DCBF constraint are imposed at each time step within a multi-step optimization model, it would result in a substantial computational burden. To reduce this complexity, we roll out the time $k$ and replace the RHS of each DCBF constraint with $h_{i}(\mathbf{x}_{t})$, as follows:
@@
Replacing (23e) with (24) can accelerate the computation. This modification affects neither the system feasibility nor its safety.